Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-05-17
ISS On-Orbit Status 05/17/08

Saturday - light duty day for CDR Volkov, FE-1 Kononenko and FE-2 Reisman.

Due to last night's extended duty after the Progress 29P arrival (5:39pm EDT), the crew had an additional 3.5 hrs sleep time this morning, i.e., wakeup - 5:30am EDT; sleep time tonight is back at the regular 5:30pm.

For the ongoing US Sleep study, currently on his voluntary 'job jar' task list, FE-2 Garrett Reisman downloaded the SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) experiment data after wakeup and before breakfast from his Actiwatch to the HRF-1 (Human Research Facility 1) laptop. (To monitor his sleep/wake patterns and light exposure, the Flight Engineer wears a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout this week, for the last time. The log entries are done within 15 minutes of final awakening for seven consecutive days.)

After opening the two hatches between FGB and 29P last night (~8:20pm), Volkov & Kononenko first installed the QD (quick disconnect) screw clamps (BZV) of the docking & internal transfer mechanism (SSVP) to rigidize the coupling and assembled the ventilation/heating air duct.

Next, Kononenko performed the standard air sampling inside the Progress with the Russian AK-1M air sampler, then deactivated the cargo ship.

Much of today's work by Oleg and Sergey focused on Progress M-64/29P unloading, starting with the cosmonauts dismantling the docking mechanism (StM, Stykovochnovo mekhanizma) between the cargo ship and the FGB. (The StM is the "classic" probe-and-cone type, consisting of an active docking assembly (ASA) with a probe (SSh), which fits into the cone (SK) on the passive docking assembly (PSA) for initial soft dock and subsequent retraction to hard dock. The ASA is mounted on the Progress' cargo module (GrO), while the PSA sits on the docking ports of the SM, FGB and DC1.)

Working in the TKG cargo ship, Volkov installed the LKT local temperature sensor commutator (TA251MB) of the BITS2-12 onboard telemetry measurement system, along with its ROM unit (read-only memory, TA765B), a 1-hr. job. The LKT was subsequently switched on by the ground to complete the basic configuration. (At a later time, Sergey will complete the electronic integration of 29P into the ISS by installing the standard US-21 matching unit, another 1-hr. task. The US-21 matching unit connects the SM with the Progress motion control and DPO thrusters systems, so that they can be commanded by the SM computer system (BVS). After bolting the box down, the CDR will hook up its the telemetry (TM) connector to the BITS2-12 onboard TM system on Go from TsUP, after Moscow had inhibited data output to the VD-SU control system mode, powered off the BITS and deactivated the SKV-2 air conditioner. These systems will subsequently be turned back on. Afterwards, the Progress thrusters will be test fired to insure their functionality in providing attitude control for ISS. 29P is then fully incorporated into the steering logic for the ISS.)

FE-1 Kononenko attended to the periodic servicing of the active Russian BMP (Harmful Impurities Removal System) by starting the "bake-out" cycle to vacuum on absorbent bed #1 of the regenerable dual-channel filtration system. The regen process will be terminated tomorrow, followed by regeneration of bed #2. (Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. In order to assist in atmosphere scrubbing after the Freon-218 (Khladon) spill from the SKV-2 air conditioner on 4/24, the BMP's regeneration cycle was moded to 5 days instead of the regular 20 days.)

Having reached Day 60 of his flight, Garrett Reisman began his second session with the NASA/JSC experiment NUTRITION w/Repository (the FD30 session having been his first), for which he had to forego exercising and food intake for eight hours. Today's protocol consisted of two blood draws (for Serum & Heparin). Later, the FE-2 set up the equipment for the 24-hour urine collections which start with the first void early tomorrow morning and continue through Sunday morning. (After performing self-phlebotomy, i.e., drawing blood samples (from an arm vein), the samples were first allowed to coagulate in the Repository for 20-30 minutes, then spun in the HRF RC (Human Research Facility/Refrigerated Centrifuge) and finally placed in MELFI (Minus-Eighty Laboratory Freezer for ISS). No thruster activity was allowed during the blood drawing. The RC was later powered off after a temperature reset to limit wear on the compressor, and cleaned. Background: NUTRITION is the most comprehensive in-flight study done by NASA to date of human physiologic changes during long-duration space flight; this includes measures of bone metabolism, oxidative damage, nutritional assessments, and hormonal changes. The Clinical Nutritional Assessment profile currently required on all U.S. Astronauts collects blood and urine samples preflight and postflight. NUTRITION expands this protocol by also capturing inflight samples and an additional postflight sample. Furthermore, additional measurements are included for samples from all sessions, including additional markers of bone metabolism, vitamin status, and hormone and oxidative stressor tests. The results will be used to better understand the impact of countermeasures (exercise and pharmaceuticals) on nutritional status and nutrient requirements. The Clinical Nutritional Assessment profile (MR016L), first started on two Mir crewmembers and then on all ISS US crews, nominally consists of two pre-flight and one post-flight analysis of nutritional status, as well as an in-flight assessment of dietary intake using the FFQ (Food Frequency Questionnaire). The current NUTRITION project expands MR016L testing in three ways: Addition of in-flight blood & urine collection (made possible by MELFI), normative markers of nutritional assessment, and a return session plus 30-day (R+30) session to allow evaluation of post-flight nutrition and implications for rehabilitation.)

Volkov & Kononenko had an hour set aside for transfers from the Progress, going by an uplinked Cargo Transfer List of individual items and their intended stowage locations.

Special attention was given by the FE-1 to the transfer, setup & situational photography of the new BIO-12 REGENERATSIYA (regeneration) payload and its ULITKA (snail) container, designed to study how zero gravity impacts structural and functional recovery of damaged organs and tissues in animal tests.

The FE-2 meanwhile started the downloading of IWIS (Internal Wireless Instrumentation System) structural dynamics data recorded during the docking activities. When completed, Garrett dismantled and stowed the IWIS hardware. (Structural dynamics data of the docking were also captured by the three RSUs of the EWIS (External Wireless Instrumentation System).)

Reisman performed the periodic offloading of the Lab CCAA (Common Cabin Air Assembly) dehumidifier's condensate tank, filling a CWC (Contingency Water Container, #1062) with the collected water slated for processing. No samples were required. (Transferred quantity is determined by allowing tank and CWC equalize with each other. Tank quantity stabilizes at a value higher than the neutral point.)

Sergey completed the routine maintenance of the SM's SOZh/ECLSS system, including ASU toilet facilities systems/replaceables. (Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, and processing U.S. condensate water as it becomes available in a filled CWC from the Lab humidifier.)

The crew conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), and RED resistive exercise device (FE-2).

Afterwards, Volkov transferred the crew's exercise data file to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Correction: The STS-124/1J EVA preparations in the Airlock (A/L) scheduled yesterday for Reisman, erroneously reported here as completed, were reportedly not done but deferred.

Weekly Science Update (Expedition Seventeen -- Week 4)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module. Inc16 science data returned on 15S have been handed over to Science Team.

ANITA (Analyzing Interferometer for Ambient Air): Continuing.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): B3 & B4 plants are growing. B3 growth is slower than estimated. Some of the B1 plants are still alive.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed. Inc16 science data returned on 15S and have been handed over to Science Team.

EuTEF (European Technology Exposure Facility): The experiment platform continues to operate nominally. A DHPU (Data Handling Processing Unit) software patch is currently developed to fix the link error issues encountered with DEBIE-2 and FIPEX instruments. - DEBIE-2: Link error still under investigation; - DOSTEL: On-going science acquisition; - EuTEMP: Currently inactive as planned; - EVC: Further troubleshooting is on-going; Still problems to get high-rate data downlink; - EXPOSE: On-going science acquisition; - FIPEX: A new script was uplinked on 5/7 and terminated nominally on 5/11; a new script has been uplinked on 5/14 but not started yet; - MEDET: On-going science acquisition, running nominal; some commanding activities to adjust parameters took place on 5/15; - PLEGPAY: Instrument is powered on, but not in science acquisition mode. New Langmuir probe runs will be started in conjunction with Progress 29P docking on 5/16; - TRIBOLAB: TRIBOLAB Pin On Disk#2 (POD#2) ran nominally until 5/9. The instrument has been put in Thermal Stabilization Mode (TSM) to prepare for an attempt to restart the Pin On Disk#1 (POD#1) run that took place on 5/9. Unfortunately, the POD#1 run went immediately to Non-Nominal Mode (NNM), for the same reason as before (friction coefficient too high); Science team is now planning to start the next POD#3 run on 5/17.

FSL (Fluid Science Laboratory): FSL MIL Bus cable repair was successfully performed on 5/5. After FSL Rack Activation from ground on 5/8, the ground confirmed that the FSL cables repair activities were successful. Further troubleshooting activities will be required with CEM-U (Upper) Optical Module.

GEOFLOW: Further troubleshooting activities with CEM-U (Upper) Optical Module are required. A new lamp set is manifested on STS-124/1J. GEOFLOW start is pending further FSL troubleshooting, and remaining commissioning activities (check of FSL optical modes). This will not occur prior to 1J flight.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.


PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Inc16 samples have been downloaded with 15S and handed over to science team.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Garrett, thanks for completing your Actiwatch download/initialization activity. You are also task listed to complete your third week of required Sleep logging this week. Any additional Sleep logging is above and beyond and greatly appreciated by the PI. After completing the Sleep log, you will have one more Actiwatch download session, and then the Actiwatch doff on the Shuttle.'

SOLAR (Solar Monitoring Observatory): During the last Sun observation period, the SOLAR platform, SOVIM and SOLSPEC Sun sensors' positions have been correlated during a so-called 'criss-cross' procedure manoeuvre on 5/5 and 5/8. The last Sun observation period has closed on 5/10. - SOVIM: Waiting for the Sun - some glitches observed in the instrument telemetry during last week; - SOLSPEC: Waiting for the Sun; - SOLACES: Waiting for the Sun, some recurrent synchronization problems between its microcontrollers during last week.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The 4 ECs (Experiment Containers) of Rotor B are yet to be retrieved by the crew from the blocked Rotor B. The first troubleshooting activity, which reinserts 6 RECs (Reference ECs) into Rotor A, has been performed on 5/15. Unfortunately, while activating BIOLAB from ground after this crew activity, a caution signal was encountered. Engineering teams are analyzing the telemetry.

CEO (Crew Earth Observations): Through 5/14 the ground has received a total of 1,383 of ISS/CEO images for review and cataloging. 'We are currently reviewing imagery with camera times corresponding to the following target list requests: Galeras Volcano, Colombia (preliminarily not acquired); Lake Nasser, Egypt (excellent coverage, this target can be pulled as completed); and Vredefort Impact Crater, South Africa (preliminarily acquired with great coverage). Nice work! Your efforts to acquire imagery of Cologne, Germany have netted partly cloudy views of several cities along the Rhine: Karlsruhe, Mainz, and Koblenz, all too far south. One of your Cologne attempts captured some striking views of the world-famous vineyards along the Moselle River about 50 miles southwest of Koblenz, Germany. One of these frames will be published on NASA/GSFC's Earth Observatory website this weekend. Good eye! Clearly you are making a most diligent effort to respond to our target requests (nearly 400 frames so far). Thank you!'

CEO (Crew Earth Observations) photo target uplinked for today were High Central Andean Glaciers, S. America (weather was favorable for photography of the summit glaciers of the Peruvian Andes. Looking to the left of track as ISS orbit paralleled the South American coastline. Imagery of the upper flanks and summits of the mountains was requested in order to track changes in ice extent and snow cover.

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use